# Twist and higher modes of a complex scalar field at the threshold of collapse

### v1

@@ I Introduction
in standard cylindrical polar coordinates. As a matter of fact, introducing angular momentum to the spacetime with scalar field matter content yields neither a spherically symmetric nor an obvious vacuum limit case that could serve as a basis for comparison of the threshold solutions. More precisely, no spherical limit exists in this case, in the sense that even perturbative solutions within our symmetry class are not spherical. Likewise, in vacuum axisymmetry, black holes with angular momentum can not be formed from regular initial data (see App. A for details). In contrast to the $m=0$ case [2], when adding net angular momentum in this manner and fixing the azimuthal winding number $m=1$, it was found in [1] that the threshold of collapse is again rendered universal. This part of the solution space is a particularly interesting, since it is not spherical yet displays critical behavior familiar from that setting. Studying it in detail may therefore strengthen our understanding of critical phenomena and shed light on the loss of universality at the axisymmetric threshold of collapse. Presently we look at different families of initial data, extending to the higher angular mode $m=2$.
@@ III Formulation and numerical methods
In order to implement this system numerically in axisymmetry we have developed a generalized symmetry reduction method in bamps, adapted to the symmetries of the problem. The bamps code is a pseudospectral code for the time development of first-order symmetric hyperbolic systems. The name chosen for the new symmetry reduction method is m-cartoon method which is a generalization of the cartoon method, simply adapted to spacetimes where there is a net angular momentum in the complex scalar field content. In this section we give a brief summary of the continuum equations solved, our grid setup, initial data solver, $hp$-mesh-refinement strategy and our improved constraint damping parameters. Full details regarding these aspects of the code can be found in the technical references [21, 22, 23, 24]. Following these, a detailed explanation of the novel m-cartoon method is provided. We end the section with an overview of the adjustments we made to our apparent horizon finder to include twist.
@@ III.1.2 Matter
I 1 1 tuned 0.6 1.0 1.0 -1 $a^{\textrm{Re}}$ 0.6 1.0 1.0 0.1
II 1 1 $3a^{\textrm{Im}}$ 0.6 1.0 1.0 -1 tuned 0.6 1.0 1.0 0.1
III 2 1 tuned 0.6 1.0 1.0 -1 $a^{\textrm{Re}}$ 0.6 1.0 1.0 0.1
IV 2 1 tuned 0.8 1.0 1.0 -1 $2a^{\textrm{Re}}$ 0.8 1.0 1.0 0.2
@@ III.1.3 Initial data
The initial data containing angular momentum follows the form given in [1], adapted to our evolved variables. However, the main difference is that in our case, a symmetric pulse about the origin has been added to ensure regularity at the center, since our pseudospectral code absolutely requires smooth data to perform accurately. The initial data of [1] are not regular, which renders their reproduction impossible using pseudospectral methods without more care. The analytic expressions we use for our initial data solver are as follows
@@
for the imaginary part.
@@ III.4 The m-cartoon method
Working to restrict to the $y=0$ plane now, we use the fact that
@@
[…]
where the derivative terms on the right-hand-side can again be replaced by the respective spatial reduction variables. Taking now the limit of $x\to 0$ for each of the components leads to
@@
$\lim_{x\to 0}\partial_{y}\Phi_{x}$ $\simeq\frac{1}{2x}\left(\partial_{x}\psi+x\partial^{2}_{x}\psi-\partial_{x}\psi\right)$
$\simeq\frac{1}{2}\partial^{2}_{x}\psi\simeq\frac{1}{2}\partial_{x}\Phi_{x}$
@@
$\lim_{x\to 0}\partial_{y}\Phi_{y}$ $\simeq\frac{1}{2x}\left(\partial_{x}\psi+x\partial^{2}_{x}\psi-m^{2}\partial_{x}\psi\right)$
$\simeq\frac{1}{2}\left(\partial^{2}_{x}\psi+\partial^{2}_{x}\psi+x\partial^{3}_{x}\psi-m^{2}\partial^{2}_{x}\psi\right)$
$\simeq\left(1-\frac{1}{2}m^{2}\right)\partial^{2}_{x}\psi\simeq\left(1-\frac{1}{2}m^{2}\right)\partial_{x}\Phi_{x}$
@@
$\lim_{x\to 0}\partial_{y}\Phi_{z}$ $\simeq im\partial_{x}\Phi_{z}\,$, (36)
where again first derivatives of $\psi$ can be replaced with the appropriate reduction variable. These relations give us a prescription for the calculation of the $y$-derivatives of the spatial reduction variables of the complex scalar field on the $y=0$ plain and in the limit of the $x=0$ axis, meaning the $z$-axis. The way this information is passed into the code is by splitting the expressions into real and imaginary parts, since the code solves for them separately. A final detail required for the numerical implementation concerns the parity conditions at the origin: along the $x$-direction, the field exhibits even parity for even $m$ and odd parity for odd $m$.
@@ III.5 Classification of spacetimes with twist
For a comprehensive review of horizon-finding techniques in numerical relativity, see [39]. Examples of horizon searches beyond the star-shaped assumption can be found in [40, 41]. More recent overviews are available in the form of PhD theses [42, 41].
@@ IV Numerical Results
As discussed above, the spacetimes of [1] cannot be replicated within our framework due to the requirement of smoothness with our pseudospectral approach. Nevertheless, the slight differences in our families of initial data allow us to examine universality in the $m=1$ mode. At the origin the field itself is zero, therefore in order to access the leading order non-zero part of the field we plot the first derivative in the $x$-direction normalized with the proper radius at the origin, $\rho_{c}=\sqrt{g_{\phi\phi}}$ and rescaled with $\left(\tau^{*}-\tau\right)$, as defined below. Figure 1 shows the highest subcritical evolution of families I and II from Table 1 with respect to similarity adapted time, as defined in Eq (3). The curves we obtain differ slightly from those reported in [1], with the only distinction being in the amplitudes of the real and imaginary components. A universal curve that would be the same for all families across different works would be the Kretschmann invariant at the origin or any non-zero gauge independent quantity extracted from the stress-energy tensor, such as $T_{ab}T^{ab}$.
Figure 1: Scaled derivatives of the scalar field at the center for the best tuned subcritical evolutions of Family I (top) and Family II (bottom). The quantity $(\tau_{*}-\tau)\,\partial_{x}\Phi/\rho_{c}$ is plotted against slow time $\ln(\tau_{*}-\tau)$. Both the real part (red dashed) and the imaginary part (blue dash-dotted) display oscillatory behavior, with the real component showing larger amplitude oscillations. While the oscillation period seems to be universal across families, the curves themselves are not. Looking at curvature scalars the spacetime metric itself is universal at the threshold.
Figure 2: Color maps of the normalized Kretschmann invariant $(\tau_{*}-\tau)\left(R_{abcd}R^{abcd}\right)^{\frac{1}{4}}$ for our best tuned subcritical runs of Family I and II initial data setups, all along the symmetry axis in single-null similarity coordinates.
@@
$\tau(t)=\int_{0}^{t}\alpha(t^{\prime},0)\,dt^{\prime}\,$, (39)
@@
$\tau_{*}=\frac{\tau_{n}\tau_{m+1}-\tau_{n+1}\tau_{m}}{\tau_{n}-\tau_{n+1}-\tau_{m}+\tau_{m+1}}\,.$ (40)
The echoing period $\Delta$ is determined using consecutive zero crossings of either the real or imaginary component of the central quantity, one obtains
@@
since each crossing interval corresponds to half a period, $\Delta/2$. The value of the period that we find for the $m=1$ families, $\Delta\simeq 0.42$, see Table 4, is consistent with the results of [1]. We note that the period $\Delta$ estimated here corresponds to the oscillation period of each individual curve representing the real and imaginary parts of the central field values, which are not themselves following universal curves (their period is universal, whereas their amplitude is not). This interpretation also applies to the results of [1], whose visualization approach we follow for comparison. In contrast, the period associated with the amplitude of the complex field or with the Kretschmann invariant at the origin would, in fact, provide a better measure of the DSS period, since those quantities exhibit universal behavior. For instance, in the $m=0$ spherical case discussed in [2], we defined as $\Delta$’ as the period of the amplitude of the complex field at the origin and obtained a value statistically identical to that of the real scalar field case. In the $m>0$ case, since the field is accessed through its rescaled spatial derivatives at the center, this procedure cannot be applied directly. Nevertheless, we confirm that the Kretschmann invariant follows a universal curve with a period of $\Delta$ for both families.
Taken together, this strongly suggests that at the threshold, the $m=1$ families of the complex scalar field coincide qualitatively with those of [1], and that there is universality and DSS. Further support comes from constructing single–null DSS–adapted coordinates, in order to show DSS behavior in an extended region, following [43, 15]. We extend the slow-time variable (now denoted $T_{\textrm{null}}$) as
@@
If the spacetime is DSS with accumulation at the center, then the coordinates $(T_{\textrm{null}},\lambda)$ are symmetry-adapted, and dimensionless scalars should appear periodic in $T_{\textrm{null}}$. We verify this by plotting heat maps of the rescaled Kretschmann scalar, $(\tau_{*}-\tau)\left(R_{abcd}R^{abcd}\right)^{\frac{1}{4}}$. Figure 2 shows the results for the best-tuned subcritical evolution family I and II. Inspecting by eye, a periodic DSS structure over an extended region is observed with period $\Delta\simeq 0.4$. We observe that in the rotating case such a null-geodesic construction remain possible along the $z$-axis, but not along the $x$-axis, where rotation causes photons to rotate around the $z$-axis.
Family m $\Delta$ $\gamma_{\textrm{sub}}$ $\gamma_{\textrm{sup}}$
I 1 $0.426\pm 0.013$ $1105\pm 0.0007 0.0$$098\pm 0.003$
II 1 $0.427\pm 0.012$ $0.1073\pm 0.0007$ $0.113\pm 0.004$
III 2 $092\pm 0.009 0.0$$349\pm 0.0003 0.0$$031\pm 0.003$
IV 2 $093\pm 0.008 0.0$$334\pm 0.0013 0.0$$0343\pm 0.0022$
Table 4: Numerical estimation of the parameters that in the Choptuik spherical spacetime appear as universal. In the first two columns the families are being specified. In the third column the period of DSS as measured by averaging over pairs of zero-crossings of the central values of the fields for each family is presented. The respective errors correspond to the standard deviations of the measured periods. In the last two columns, we show the regression results from the scaling exponents of the Ricci scalar, Eq. (45), at the subcritical side and of the apparent horizon mass, Eq. (44), at the supercritical side. The errors presented here are the errors resulting from the regression. In our study, the standard picture of universality seems to be maintained for every fixed mode $m>0$.
@@
Figure 3: Log-log plot of the maximum of the Ricci scalar, $R_{\text{max}}$, as a function of the inverse of the phase space distance to criticality, $|a-a_{*}|^{-1}$, for the $m=1$ family I. The data exhibit a clear power-law scaling, with slope $2\gamma_{\textrm{sub}}$. Fitting yields a universal critical exponent $\gamma\simeq 0.11$, in agreement with the results of [1]. The regression parameter $r^{2}$ quantifies the level of scattering of the data.
@@
is expected to be the same scaling exponent we see on the subcritical side, $\gamma_{\textrm{sup}}=\gamma_{\textrm{sub}}$. As illustrated in Fig. 4 for family I and additionally summarized in Table 4 for family II, the scaling exponent on the supercritical side of the critical regime is indeed $\gamma_{\textrm{sup}}\simeq 0.1$, demonstrating the universality of power-law scaling exponents for quantities with dimensions of length in near-threshold solutions with $m=1$.
Figure 4: Log-log plot of the mass of the first apparent horizon that appears in each simulation, $M_{\text{AH}}$, against the phase space distance to the threshold solution, $|a-a_{*}|$, for the $m=1$ family I. The mass scales with an exponent $\gamma_{\textrm{sup}}\simeq 0.1$ the same universal number found on the subcritical side.
Since the spacetime carries angular momentum, it is natural to measure it at the horizon using Eq. (38), and to study its scaling properties. This is subtle because of the foliation dependence of the apparent horizon, but is by far the most convenient option numerically. In Figure 5, we plot the apparent horizon angular momentum against the mass at the moment of formation in our near-critical $m=1$ simulations of family I. While $J_{\mathrm{AH}}$, as well as $M_{\mathrm{AH}}$, are gauge-dependent and subject to numerical noise, especially near the threshold where they become vanishingly small, yet the data nonetheless exhibit a coherent decreasing trend of $J_{\mathrm{AH}}$ as $M_{\mathrm{AH}}$ approaches the threshold. Moreover, the dimensionless spin $\chi_{\mathrm{AH}}=J_{\mathrm{AH}}/M_{\mathrm{AH}}^{2}$ systematically decreases toward the threshold, affirming that angular momentum becomes negligible in the critical regime. In our best tuned data, we have $J_{\mathrm{AH}}/M_{\mathrm{AH}}^{2}=O(10^{-4})$. By comparison, in [1] (Fig. 4) two separate scaling regimes were found — a steep $J\propto M^{6}$ regime very close to threshold, and a milder $J\propto M^{2.2}$ regime farther from it, suggesting an abrupt slope change in the log-log plot. In our results, though scatter is large, possibly due to gauge effects, we do not resolve such a broken slope. Instead the data are consistent with a single power-law form over our explored mass range. Nonetheless the qualitative agreement remains: angular momentum vanishes more rapidly than $M^{2}$, in line with the interpretation that it is irrelevant at the threshold within this family.
@@
We now investigate $m=2$ families in order to test the universality of the critical solution across higher angular modes. Since for this data the first derivative of the field at the origin is zero, we access the field through its second derivative, normalizing this time with $(\tau_{*}-\tau)^{2}$ and $\rho_{c}^{2}=g_{\phi\phi}$. Near-critical central values from the evolution of the real and imaginary parts, shown in Fig. 6, exhibit echoing behavior analogous to the $m=1$ case (compare with Fig. 1). In this case however the extracted echoing period is $\Delta\simeq 0.09$, noticeably smaller than the $m=1$ value, and thus signaling a dependence of $\Delta$ on the angular mode.
Figure 6: Scaled derivatives of the scalar field at the center for the best tuned subcritical evolutions in the $m=2$ families III and IV. The quantity $(\tau_{*}-\tau)\,\partial^{2}_{x}\Phi/\rho^{2}_{c}$ is plotted against slow time $T=-\ln(\tau_{*}-\tau)$. Both the real (red dashed) and imaginary (blue dash-dotted) parts display echoing behavior. The shape of the curves of the real and imaginary field components is not universal, but the oscillation period extracted from both families yields $\Delta\simeq 0.09$, significantly smaller than in the $m=1$ case. However, non-zero curvature scalars at the origin are universal.
@@
Figure 7: Color maps of the normalized Kretschmann invariant $(\tau_{*}-\tau)\left(R_{abcd}R^{abcd}\right)^{\frac{1}{4}}$ for our best tuned subcritical runs of Family III and IV initial data setups, all along the symmetry axis in single-null similarity coordinates.
Looking into the phase space on the subcritical side, we analyze the scaling of $R_{\text{max}}$ (Fig. 8) and obtain $\gamma_{\textrm{sub}}\simeq 0.035$.
Figure 8: Log-log plot of the maximum of the Ricci scalar, $R_{\text{max}}$, as a function of the distance to criticality, $|a-a_{*}|$, for the $m=2$ family III. The data follow a power law with two apparent slopes, but the one closer to criticality is taken as the relevant scaling. From this regime we extract $\gamma_{\textrm{sub}}\simeq 0.035$, a value compatible across all $m=2$ families (see Table 4) and smaller than the $m=1$ result.

### v2

@@ I Introduction
in standard cylindrical polar coordinates. As a matter of fact, introducing angular momentum to the spacetime with scalar field matter content yields neither a spherically symmetric nor an obvious vacuum limit case that could serve as a basis for comparison of the threshold solutions. More precisely, no spherical limit exists in this case, in the sense that even perturbative solutions within our symmetry class are not spherical. Likewise, in vacuum axisymmetry, black holes with angular momentum cannot be formed from regular initial data (see App. A for details). In contrast to the $m=0$ case [2], when adding net angular momentum in this manner and fixing the azimuthal winding number $m=1$, it was found in [1] that the threshold of collapse is again rendered universal, for numerous families of $l=|m|=1$ initial data. This part of the solution space is a particularly interesting, since it is not spherical yet displays critical behavior familiar from that setting. Studying it in detail may therefore strengthen our understanding of critical phenomena and shed light on the loss of universality at the axisymmetric threshold of collapse. Presently we look at different families of initial data, extending to the higher angular mode $m=2$.
@@ III Formulation and numerical methods
In order to implement this system numerically in axisymmetry we have developed a generalized symmetry reduction method in bamps, adapted to the symmetries of the problem. The bamps code is a pseudospectral code for the time development of first-order symmetric hyperbolic systems. The name chosen for the new symmetry reduction method is m-cartoon method. It is a generalization of the cartoon method, simply adapted to spacetimes where there is a net angular momentum in the complex scalar field content. In this section we give a brief summary of the continuum equations solved, our grid setup, initial data solver, $hp$-mesh-refinement strategy and our improved constraint damping parameters. Full details regarding these aspects of the code can be found in the technical references [21, 22, 23, 24]. Following these, a detailed explanation of the novel m-cartoon method is provided. We end the section with an overview of the adjustments we made to our apparent horizon finder to include twist.
@@ III.1.2 Matter
I 1 1 tuned 0.6 1.0 1.0 -1 $a^{\textrm{Re}}$ 0.6 1.0 1.0 0.1
II 1 1 $3a^{\textrm{Im}}$ 0.6 1.0 1.0 -1 tuned 0.6 1.0 1.0 0.1
III 2 1 tuned 0.6 1.0 1.0 -1 $a^{\textrm{Re}}$ 0.6 1.0 1.0 0.1
IV 2 1 tuned 0.8 1.0 1.0 -1 $2a^{\textrm{Re}}$ 0.8 1.0 1.0 0.2
@@ III.1.3 Initial data
The initial data containing angular momentum follows the form given in [1], adapted to our evolved variables, $\left\{\psi_{\textrm{Re}},\psi_{\textrm{Im}},\Pi_{\textrm{Re}},\Pi_{\textrm{Im}}\right\}$. However, the main difference is that in our case, a symmetric pulse about the origin has been added to ensure regularity at the center, since our pseudospectral code absolutely requires smooth data to perform accurately. The initial data of [1] are not regular, which renders their reproduction impossible using pseudospectral methods without more care. The analytic expressions we use for our initial data solver are as follows
@@
for the imaginary part. We observe in passing that the these scalar field data consist of pure $l=|m|$ spherical harmonic modes only. This property is not likely to be propagated in strong-field time evolutions.
@@ III.4 The m-cartoon method
To replace $\partial_{\rho}\psi$, we use the fact that
@@
$\begin{split}\partial_{y}\Phi_{i}={}&im\left(\frac{1}{x}\partial_{x}\psi,0,\frac{\partial_{z}\psi}{x}\right)+\left(0,\frac{1}{x}\partial_{x}\psi,0\right)\\ &+im\left(-\frac{1}{x^{2}},0,0\right)\psi-m^{2}\left(0,\frac{1}{x^{2}},0\right)\psi\,.\end{split}$ (35)
Taking now the limit of $x\to 0$ for each of the components leads to
@@
$\lim_{x\to 0}\partial_{y}\Phi_{x}$ $=\lim_{x\to 0}\frac{im}{2x}\left(\partial_{x}\psi+x\partial^{2}_{x}\psi-\partial_{x}\psi\right)$
$=\frac{im}{2}\partial^{2}_{x}\psi=\frac{im}{2}\partial_{x}\Phi_{x}$
@@
$\lim_{x\to 0}\partial_{y}\Phi_{y}$ $=\lim_{x\to 0}\frac{1}{2x}\left(\partial_{x}\psi+x\partial^{2}_{x}\psi-m^{2}\partial_{x}\psi\right)$
$=\frac{1}{2}\left(\partial^{2}_{x}\psi+\partial^{2}_{x}\psi+x\partial^{3}_{x}\psi-m^{2}\partial^{2}_{x}\psi\right)$
$=\left(1-\frac{1}{2}m^{2}\right)\partial^{2}_{x}\psi=\left(1-\frac{1}{2}m^{2}\right)\partial_{x}\Phi_{x}$
@@
$\lim_{x\to 0}\partial_{y}\Phi_{z}$ $=im\partial_{x}\partial_{z}\psi=im\partial_{x}\Phi_{z}\,$, (36)
where second derivatives of $\psi$ have been replaced using the appropriate reduction variables in the code. These relations give us a prescription for the calculation of the $y$-derivatives of the spatial reduction variables of the complex scalar field on the $y=0$ plain and in the limit of the $x=0$ axis, meaning the $z$-axis. The way this information is passed into the code is by splitting the expressions into real and imaginary parts, since the code solves for them separately. A final detail required for the numerical implementation concerns the parity conditions at the origin: along the $x$-direction, the field exhibits even parity for even $m$ and odd parity for odd $m$.
@@ III.5 Classification of spacetimes with twist
For a comprehensive review of horizon-finding techniques in numerical relativity, see [39]. Examples of horizon searches beyond the star-shaped assumption can be found in [40]. More recent overviews are available in the form of PhD theses [41, 42].
@@ IV Numerical Results
As discussed above, the spacetimes of [1] cannot be replicated within our framework due to the requirement of smoothness with our pseudospectral approach. Incidentally, the degree of smoothness is expected in general to effect the behavior of near the threshold. We are concerned with the category of smooth solutions. In any case, the slight differences in our families of initial data allow us to examine universality in the $m=1$ mode. At the origin the field itself is zero, therefore in order to access the leading order non-zero part of the field we plot the first derivative $\partial_{{\rho_{c}}}\psi$ with the proper radius at the origin, $\rho_{c}=\sqrt{g_{\phi\phi}}\stackrel{{\scriptstyle y=0}}{{=}}\rho\sqrt{g_{yy}}=x\sqrt{g_{yy}}$ and rescaled with $\left(\tau^{*}-\tau\right)$, as defined in the introduction. Figure 1 shows the highest subcritical evolution of families I and II from Table 1 with respect to similarity adapted time, as defined in Eq (3). The curves we obtain differ slightly from those reported in [1], with the only distinction being in the amplitudes of the real and imaginary components. A universal curve that would be the same for all families across different works would be the Kretschmann invariant at the origin or any non-zero gauge independent quantity extracted from the stress-energy tensor, such as $T_{ab}T^{ab}$.
Figure 1: Scaled derivatives of the scalar field at the center for the best tuned subcritical evolutions of Family I (top) and Family II (bottom). The quantity $(\tau_{*}-\tau)\,\partial_{\rho_{c}}\psi$ is plotted against slow time $\ln(\tau_{*}-\tau)$. Both the real part (red dashed) and the imaginary part (blue dash-dotted) display oscillatory behavior, with the real component showing larger amplitude oscillations. While the oscillation period seems to be universal across families, the curves themselves are not. Looking at curvature scalars the spacetime metric itself is universal at the threshold.
Figure 2: Color maps of the normalized Kretschmann invariant $(\tau_{*}-\tau)\left|R_{abcd}R^{abcd}\right|^{\frac{1}{4}}$ for our best-tuned subcritical runs of Family I and II initial data setups, all along the symmetry axis in single-null similarity coordinates.
@@
$\tau(t)=\int_{0}^{t}\alpha(t^{\prime},x^{i}=0)\,dt^{\prime}\,$, (39)
@@
$\tau_{*}=\frac{\tau_{n}\tau_{m+1}-\tau_{n+1}\tau_{m}}{\tau_{n}-\tau_{n+1}-\tau_{m}+\tau_{m+1}}\,$, (40)
and, subsequently, average over multiple pairs of zero crossings. The echoing period $\Delta$ is determined using consecutive zero crossings of either the real or imaginary component of the central quantity, one obtains
@@
since each crossing interval corresponds to half a period, $\Delta/2$. Again, we average over multiple consecutive zero crossings. The value of the period that we find for the $m=1$ families, $\Delta\simeq 0.42$, see Table 4, is consistent with the results of [1]. We note that the period $\Delta$ estimated here corresponds to the oscillation period of each individual curve representing the real and imaginary parts of the central field values, which are not themselves following universal curves (their period is universal, whereas their amplitude is not.
The Kretschmann invariant on the other hand provides better measure of the DSS period, because it exhibits universal behavior. For instance, in the $m=0$ spherical case discussed in [2], the period of the Kretschmann invariant was statistically identical to that of the real scalar field. We confirm that the central value of the Kretschmann invariant against slow time at the origin also follows a universal curve with a period of $\Delta_{I}\simeq 0.43\pm 0.03$ for both of our $m=1$ families.
Taken together, this strongly suggests that at the threshold, the $m=1$ families of the complex scalar field coincide qualitatively with those of [1], and that there is universality and DSS. Further support comes from constructing single-null DSS-adapted coordinates, in order to show DSS behavior in an extended region, following [43, 15]. We extend the slow-time variable (now denoted $T_{\textrm{null}}$) as
@@
If the spacetime is DSS with accumulation at the center, then the coordinates $(T_{\textrm{null}},\lambda)$ are symmetry-adapted, and dimensionless scalars should appear periodic in $T_{\textrm{null}}$. We verify this by plotting heat maps of the rescaled Kretschmann scalar, $(\tau_{*}-\tau)\left|R_{abcd}R^{abcd}\right|^{\frac{1}{4}}$. Figure 2 shows the results for the best-tuned subcritical evolution family I and II. Inspecting by eye, a periodic DSS structure over an extended region away from the origin is observed with period $\Delta\simeq 0.4$, compatible with the scalar field result. We observe that in the rotating case such a null-geodesic construction remain possible along the $z$-axis, but not along the $x$-axis, where rotation causes photons to rotate around the $z$-axis.
Family m $\Delta$ $\Delta_{I}$ $\gamma_{\textrm{sub}}$ $\gamma_{\textrm{sup}}$
I 1 $0.426\pm 0.013$ $0.43\pm 0.03$ $1105\pm 0.0007 0.0$$098\pm 0.003$
II 1 $0.427\pm 0.012$ $0.429\pm 0.019$ $0.1073\pm 0.0007$ $0.113\pm 0.004$
III 2 $092\pm 0.009 0.0$$92\pm 0.007 0.0$$349\pm 0.0003 0.0$$031\pm 0.003$
IV 2 $093\pm 0.008 0.0$$88\pm 0.004 0.0$$334\pm 0.0013 0.0$$0343\pm 0.0022$
Table 4: Numerical estimation of the parameters that in the Choptuik spherical spacetime appear as universal. In the first two columns the families are being specified. In the third column the period of DSS as measured by averaging over pairs of zero-crossings of the central values of the fields for each family is presented. The respective errors correspond to the standard deviations of the measured periods. In the fourth column, we calculate the period, $\Delta_{I}$, of the central value of the Kretschmann invariant, II, with respect to slow time, using the maxima of the curve. In the last two columns, we show the regression results from the scaling exponents of the Ricci scalar, Eq. (48), at the subcritical side and of the apparent horizon mass, Eq. (44), at the supercritical side. The errors presented here are the errors resulting from the regression. In our study, the standard picture of universality seems to be maintained for every fixed mode $m>0$.
@@
Figure 3: Global maximum of the Ricci scalar, $R_{\text{max}}$, as a function of the inverse of the phase space distance to criticality, $|a-a_{*}|^{-1}$, for the $m=1$ family I. The data exhibit a clear power-law scaling, with slope $2\gamma_{\textrm{sub}}$. Fitting yields a universal critical exponent $\gamma\simeq 0.11$, in agreement with the results of [1]. The regression parameter $r^{2}$ quantifies the level of scattering of the data.
@@
is expected to be the same scaling exponent we see on the subcritical side, $\gamma_{\textrm{sup}}=\gamma_{\textrm{sub}}$. As illustrated in Fig. 4 for family I and additionally summarized in Table 4 for family II, the scaling exponent on the supercritical side of the critical regime is indeed $\gamma_{\textrm{sup}}\simeq 0.1$, demonstrating the universality of power-law scaling exponents for quantities with dimensions of length in near-threshold solutions with $m=1$. Upon closer inspection the error margins reported in Table 4 seem to indicate a discrepancy between the super- and subcritical scaling exponents, but this can be explained by the fact that the error margins do not account for the systematic error due to finite resolution of the simulation and the foliation dependence of the apparent horizon.
Figure 4: Mass of the first apparent horizon that appears in each simulation, $M_{\text{AH}}$, against the phase space distance to the threshold solution, $|a-a_{*}|$, for the $m=1$ family I. The mass scales with an exponent $\gamma_{\textrm{sup}}\simeq 0.1$ the same universal number found on the subcritical side.
Since the spacetime carries angular momentum, it is natural to measure it at the horizon using Eq. (38), and to study its scaling properties. This is subtle because of the foliation dependence of the apparent horizon, but is by far the most convenient option numerically. In Fig. 5, we plot the apparent horizon angular momentum against the mass at the moment of formation in our near-critical $m=1$ simulations of family I. While $J_{\mathrm{AH}}$, as well as $M_{\mathrm{AH}}$, are gauge-dependent and subject to numerical noise, especially near the threshold where they become vanishingly small, yet the data nonetheless exhibit a coherent decreasing trend of $J_{\mathrm{AH}}$ as $M_{\mathrm{AH}}$ approaches the threshold. Moreover, the dimensionless spin $\chi_{\mathrm{AH}}=J_{\mathrm{AH}}/M_{\mathrm{AH}}^{2}$ systematically decreases toward the threshold, affirming that angular momentum becomes negligible in the critical regime. In our best-tuned data, we have the dimensionless spin is on the order of $10^{-4}$. From a fit to the data we obtain a scaling of $J_{\mathrm{AH}}~\propto M_{\mathrm{AH}}^{2.73}$, as shown in Fig. 5. By comparison, in [1] (see Fig. 4 of that paper) two separate scaling regimes were found – a steep $J\propto M^{6}$ regime very close to threshold, and a milder $J\propto M^{2.2}$ regime farther from it, suggesting an abrupt slope change in the log-log plot. In our results, though scatter is large, possibly due to gauge effects, we do not resolve such a broken slope. Instead the data are consistent with a single power-law form over our explored mass range. Nonetheless the qualitative agreement remains: angular momentum vanishes more rapidly than $M^{2}$, in line with the interpretation that it is irrelevant at the threshold within this family.
@@
We now investigate $m=2$ families in order to test the universality of the critical solution across higher angular modes. Since for this data the first derivative of the field at the origin is zero, we access the field through its second derivative, $\partial^{2}_{\rho_{c}}\psi$, with $\rho_{c}=\sqrt{g_{\phi\phi}}$ being the proper radius at the origin, normalizing this time with $(\tau_{*}-\tau)^{2}$. Near-critical central values from the evolution of the real and imaginary parts, shown in Fig. 6, exhibit echoing behavior analogous to the $m=1$ case (compare with Fig. 1). In this case however the extracted echoing period is $\Delta\simeq 0.09\pm 0.01$, noticeably smaller than the $m=1$ value, and thus signaling a dependence of $\Delta$ on the angular mode. We confirm that the central value of the Kretschmann invariant against slow time at the origin also follows a universal curve with a period of $\Delta_{I}\simeq 0.09\pm 0.01$ for both of our $m=1$ families.
Figure 6: Scaled derivatives of the scalar field at the center for the best-tuned subcritical evolutions in the $m=2$ families III and IV. The quantity $(\tau_{*}-\tau)\,\partial^{2}_{\rho_{c}}\psi$ is plotted against slow time $T=-\ln(\tau_{*}-\tau)$. Both the real (red dashed) and imaginary (blue dash-dotted) parts display echoing behavior. The shape of the curves of the real and imaginary field components is not universal, but the oscillation period extracted from both families yields $\Delta\simeq 0.09$, significantly smaller than in the $m=1$ case. However, non-zero curvature scalars at the origin are universal.
@@
Figure 7: Color maps of the normalized Kretschmann invariant $(\tau_{*}-\tau)\left|R_{abcd}R^{abcd}\right|^{\frac{1}{4}}$ for our best-tuned subcritical runs of Family III and IV initial data setups, all along the symmetry axis in single-null similarity coordinates.
Looking into the phase space on the subcritical side, we analyze the scaling of $R_{\text{max}}$ (Fig. 8) of family III and obtain $\gamma_{\textrm{sub}}\simeq 0.035$.
Figure 8: Global maximum of the Ricci scalar, $R_{\text{max}}$, as a function of the distance to criticality, $|a-a_{*}|$, for the $m=2$ family III. The data follow a power law with two apparent slopes, but the one closer to criticality is taken as the relevant scaling. From this regime we extract $\gamma_{\textrm{sub}}\simeq 0.035$, a value compatible across all $m=2$ families (see Table 4) and smaller than the $m=1$ result.